# Foundation Policies with Hilbert Representations

### v1

@@ Abstract
Unsupervised and self-supervised objectives, such as next token prediction, have enabled pre-training generalist models from large amounts of unlabeled data. In reinforcement learning (RL), however, finding a truly general and scalable unsupervised pre-training objective for generalist policies from offline data remains a major open question. While a number of methods have been proposed to enable generic self-supervised RL, based on principles such as goal-conditioned RL, behavioral cloning, and unsupervised skill learning, such methods remain limited in terms of either the diversity of the discovered behaviors, the need for high-quality demonstration data, or the lack of a clear prompting or adaptation mechanism for downstream tasks. In this work, we propose a novel unsupervised framework to pre-train generalist policies that capture diverse, optimal, long-horizon behaviors from unlabeled offline data such that they can be quickly adapted to any arbitrary new tasks in a zero-shot manner. Our key insight is to learn a structured representation that preserves the temporal structure of the underlying environment, and then to span this learned latent space with directional movements, which enables various zero-shot policy “prompting” schemes for downstream tasks. Through our experiments on simulated robotic locomotion and manipulation benchmarks, we show that our unsupervised policies can solve goal-conditioned and general RL tasks in a zero-shot fashion, even often outperforming prior methods designed specifically for each setting. Our code and videos are available at https://seohong.me/projects/hilp/.
@@ 3 Preliminaries and Problem Setting
Hilbert space. A Hilbert space ${\mathcal{Z}}$ is a complete vector space equipped with an inner product $\langle x,y\rangle$, the induced norm $\|x\|=\sqrt{\langle x,x\rangle}$, and the induced metric $d(x,y)=\|x-y\|$ for $x,y\in{\mathcal{Z}}$. Intuitively, a Hilbert space can roughly be thought of as a “stricter” version of a metric space, where there exists an inner product that is consistent with the metric. For example, a Euclidean space with the $\ell^{1}$- or $\ell^{\infty}$-norm is a metric space but not a Hilbert space, whereas a Euclidean space with the $\ell^{2}$-norm is a Hilbert space, as $\|x\|_{2}=\sqrt{x^{\top}x}$ for $x\in{\mathbb{R}}^{D}$. In our experiments, we will mainly employ Euclidean spaces (with the $\ell^{2}$-norm) as Hilbert spaces, but the theorems in the paper can be applied to any arbitrary Hilbert space.
@@
Our goal is to pre-train a versatile latent-conditioned policy $\pi(a\mid s,z)$, where $z\in{\mathcal{Z}}$ denotes a latent vector (which we call a task or a skill), purely from the unlabeled offline data ${\mathcal{D}}$, without online interactions. For the evaluation of the pre-trained policy, we consider three evaluation settings. (1) Zero-shot RL: Given a reward function $r(s)$, we aim to find the best latent vector $z$ that maximizes the reward function, without additional training. (2) Zero-shot goal-conditioned RL: Given a target goal $g\in{\mathcal{S}}$, we aim to find the best latent vector $z$ of the policy $\pi(a\mid s,z)$ that leads to the goal as quickly as possible, without additional training. The goal is specified at test time. (3) Hierarchical RL: Given a reward function $r(s)$, we train a high-level policy $\pi^{h}(z\mid s)$ that sequentially combines pre-trained skills to maximize the reward function using offline RL. In all three settings, we only allow online interaction with the environment during the final evaluation.
@@ 4.1 Hilbert Representations
Remarks. There exist three potential limitations with Equation 6. First, the distance metric in ${\mathcal{Z}}$ is symmetric, whereas temporal distances might be asymmetric. Second, even when the environment dynamics are symmetric, there might not exist an exact isometry between the MDP and the Hilbert space (Indyk et al., 2017; Pitis et al., 2020). Third, we use a discount factor $\gamma$ in Equation 6, but temporal distances are undiscounted. In this regard, our objective might better be viewed as finding the best discounted Hilbert approximation of the MDP, rather than learning an exact Hilbert abstraction. While this might not be ideal in highly asymmetric environments, we note that we will not directly use the potentially erroneous parameterized value function $V(s,g)$ for policy learning. We will instead only take the representation $\phi$, defining a new reward function as well as a new value function to pre-train an unsupervised latent-conditioned policy, as we will describe in Section 4.2. After all, our goal is to train a representation $\phi$ that captures the long-term temporal structure of the MDP, and we empirically found that even approximate Hilbert representations lead to diverse useful behaviors that can be directly used to solve downstream tasks in our experiments (Section 6). We refer to Appendix C for further discussion.
@@ 5 Why are HILPs useful?
HILPs, in combination with structured representations $\phi(s)$, provide a number of ways to solve downstream tasks efficiently, often even in a zero-shot fashion. In this section, we describe three test-time “prompting” strategies for HILPs for zero-shot RL (Section 5.1), zero-shot goal-conditioned RL (Section 5.2), and test-time planning (Section 5.3).
@@ 5.1 Zero-Shot RL
5.2 Zero-Shot Goal-Conditioned RL
@@ 6 Experiments
In our experiments, we empirically evaluate the performance of HILPs on various types of downstream tasks. In particular, we consider three different experimental settings: zero-shot RL, zero-shot goal-conditioned RL, and hierarchical RL (see Section 3 for the problem statements), in which we aim to answer the following questions: (1) Can HILPs be prompted to solve goal-conditioned and reward-based tasks in a zero-shot manner? (2) Can a single general HILP framework outperform prior state-of-the-art approaches specialized in each setting? (3) Does test-time planning improve performance? In our experiments, we use $8$ random seeds unless otherwise stated, and report $95\%$ bootstrap confidence intervals in plots and standard deviations in tables. Our code is available at this repository.
@@ 6.1 Zero-Shot RL
6.2 Zero-Shot Goal-Conditioned RL
@@
Table 1: Zero-shot goal-conditioned RL performance ($\mathbf{8}$ seeds). HILP achieves the best performance among general offline unsupervised policy learning methods, while being comparable to methods that are specifically designed to solve offline goal-conditioned RL. With our efficient test-time planning procedure based on Hilbert representations, HILPs often even outperform offline goal-conditioned RL methods.
@@
visual-kitchen-partial $\mathbf{62.6}$$\pm 6.4$ $57.8$$\pm 7.4$ $46.3$$\pm 2.7$ $40.8$$\pm 3.2$ $57.8$$\pm 4.8$
visual-kitchen-mixed $\mathbf{57.5}$$\pm 4.2$ $52.9$$\pm 4.7$ $50.9$$\pm 3.6$ $49.6$$\pm 5.7$ $\mathbf{55.4}$$\pm 4.6$
Table 1 shows the results, which suggest HILPs can solve challenging long-horizon goal-conditioned tasks in a zero-shot manner. HILPs significantly outperform previous general unsupervised policy learning methods (FB and three SF methods), and with test-time planning, they often even outperform GCRL-dedicated methods (e.g., GC-IQL). We note that this planning procedure is efficiently done only with elementary algebraic operations based on pre-computed representations (Section 5.3), thanks to our structured Hilbert representations. To further study the effect of test-time planning, we compare the performances of HILPs with different numbers (00-$3$) of midpoint recursions. We report results in Figure 7, which shows that iterative refinement of prompts via our test-time planning approach improves performance consistently. We refer to Appendix B for further analyses, including an ablation study and latent space visualization.
@@ Appendix A Limitations
As mentioned in Section 4.1, one limitation of our Hilbert representation objective is that a symmetric Hilbert space might not be expressive enough to capture arbitrary MDPs. Although we empirically demonstrate that HILPs exhibit strong performances in various complex, long-horizon simulated robotic environments, they might struggle in highly asymmetric or disconnected MDPs, where there might not exist a reasonable approximate isometry to a Hilbert space. We believe this limitation may be resolved by combining the notion of an inner product with a universal quasimetric embedding (Pitis et al., 2020; Wang et al., 2023), which we leave for future work. Another limitation is that our value-based representation learning objective (Equation 6) might be optimistically biased in stochastic or partially observable environments (Park et al., 2023). We believe combining our method with history-conditioned policies or recurrent world models may be one possible solution to deal with such MDPs. Finally, we use Euclidean spaces as Hilbert spaces for our experiments in this work. We believe applying HILPs to more general Hilbert spaces, such as the $L^{2}$ function space or reproducing kernel Hilbert spaces, may significantly enhance the expressivity of Hilbert representations.
@@
Figure 9: Ablation study of latent dimensions for zero-shot goal-conditioned RL ($\mathbf{4}$ seeds). In D4RL tasks, $D=32$ generally leads to the best performance across the environments.
@@ Appendix B Additional Results
Ablation study. Figures 9 and 8 show how the dimension $D$ of the latent space ${\mathcal{Z}}={\mathbb{R}}^{D}$ affects the performances of zero-shot RL and zero-shot goal-conditioned RL (without planning), where we use $D=50$ for zero-shot RL and $D=32$ for goal-conditioned RL in our main experiments. The results suggest that a latent dimension between $25$ and $64$ generally leads to the best performance in both cases.
@@
Full results on the ExORL benchmark. Tables 3 and 4 present the full results (unnormalized returns) on the state- and pixel-based ExORL benchmarks. Figures 11, 13, 12, 14, 16 and 15 show plots with three different aggregation criteria (per-environment, per-dataset, and per-task) on the state- and pixel-based ExORL benchmarks.
@@ Appendix C Theoretical Results
Limitations. One natural question to ask is whether it is always possible to embed any MDP into a Hilbert space up to arbitrary accuracy. Unfortunately, this is not always possible. First, temporal distances are asymmetric but the distance metric of the Hilbert space ${\mathcal{Z}}$ is symmetric. Second, even when the environment is completely symmetric, there exists a symmetric MDP that is not embeddable into a Hilbert space with an arbitrarily low approximation error (Indyk et al., 2017; Pitis et al., 2020). This is mainly because Hilbert spaces are highly structured, especially compared to metric or quasimetric spaces, which do not require a well-defined inner product. Nonetheless, the inner product structure of the Hilbert space naturally enables useful prompting strategies for zero-shot RL and zero-shot goal-conditioned RL, and we empirically found that even MDPs that in principle do not have a lossless Hilbert representation can still be solved effectively via our method in our experiments. We believe finding a way to relax the Hilbert condition while having similar prompting and planning strategies is an exciting and important future research direction.
@@ Appendix D Experimental Details
We implement HILPs based on two different codebases: the official implementation of FB representations (Touati et al., 2023) for zero-shot RL experiments and that of HIQL (Park et al., 2023) for zero-shot goal-conditioned RL and hierarchical RL experiments. Our implementations are publicly available at the following repository: https://github.com/seohongpark/HILP. We run our experiments on an internal cluster consisting of A5000 GPUs. Each run in this work takes no more than $28$ hours.
@@ D.2 Implementation Details
Hilbert representations. In Equation 6, we use the same goal relabeling strategy as Park et al. (2023) except that we do not set $g=s$ since $V(s,s)=0$ is always guaranteed in our parameterization. Namely, we sample $g$ either from a geometric distribution over the future states within the same trajectory (with probability $0.625$), or uniformly from the dataset (with probability $0.375$). To avoid numerical issues with gradient descent, we add a small value (${\varepsilon}=10^{-6}$) when computing $\|\phi(s)-\phi(g)\|$.
@@
Zero-shot goal-conditioned RL (Section 6.2). We implement HILP on top of the official codebase of the work by Park et al. (2023). For HILP, we use the reward function in Equation 7 and the goal-conditioned prompting scheme introduced in Section 5.2. We use IQL (Kostrikov et al., 2022) with AWR (Peng et al., 2019) as an offline algorithm to train policies. For HILP-Plan, at each evaluation epoch, we first randomly sample $N=50000$ states $w_{1},w_{2},\dots,w_{N}$ from the dataset ${\mathcal{D}}$, and pre-compute their representations $\phi(w_{1}),\phi(w_{2}),\dots,\phi(w_{N})$. Then, at every time step, we find the $\operatorname*{arg\,min}$ of Equation 10 over the $N$ samples using the pre-computed representations. In practice, we use the average of the $50$ $\operatorname*{arg\,min}$ representations, as we found this to lead to better performance. For GC-IQL and GC-BC, we use the implementations provided by Park et al. (2023). They are implemented on the same codebase as HILP. For GC-CQL, we modify the JaxCQL repository (Geng, 2022) to make it compatible with our goal-conditioned setting. We mostly follow the hyperparameters used by Nakamoto et al. (2023). We use the same goal relabeling strategy as Park et al. (2023) for all three goal-conditioned RL methods. For FB, we use the official implementation provided by Touati et al. (2023), where we additionally implement D4RL environments. For SF methods (FDM, Lap, and Rand), we re-implement IQL versions of them on the same codebase as HILP, as we found these versions to perform better than the original implementations by Touati et al. (2023). Among FB and SF methods, we only re-implement SF methods based on IQL, as FB in its current form is not directly compatible with IQL. We report the full list of the hyperparameters used in our zero-shot goal-conditioned RL experiments in Table 6.
@@
MLP dimensions $(512,512)$ ($\phi$), $(1024,1024,1024)$ (others)111 Following Touati et al. (2023), for policies $\pi(a\mid s,z)$ and TD3 values $Q(s,a,z)$, we process $s$ (or $(s,a)$) and $(s,z)$ separately with $(1024,512)$-sized MLPs, concatenate them together, and then pass another $(1024)$-sized MLP.
@@
Table 6: Hyperparameters for zero-shot goal-conditioned RL.
@@
Policy MLP dimensions $(256,256)$ ({GC-BC, GC-IQL} on AntMaze-Large)222 We found that $(256,256)$-sized policy networks lead to better performance than $(512,512,512)$-sized ones for GC-BC and GC-IQL on AntMaze-Large. , $(512,512,512)$ (others)
@@
Latent dimension $32$ (HILP), $8$ (OPAL)333 We found that OPAL works better with $8$-dimensional latent spaces (as in the original work), compared to $32$-dimensional ones (as in this work), especially in AntMaze tasks.

### v2

@@ Abstract
Unsupervised and self-supervised objectives, such as next token prediction, have enabled pre-training generalist models from large amounts of unlabeled data. In reinforcement learning (RL), however, finding a truly general and scalable unsupervised pre-training objective for generalist policies from offline data remains a major open question. While a number of methods have been proposed to enable generic self-supervised RL, based on principles such as goal-conditioned RL, behavioral cloning, and unsupervised skill learning, such methods remain limited in terms of either the diversity of the discovered behaviors, the need for high-quality demonstration data, or the lack of a clear adaptation mechanism for downstream tasks. In this work, we propose a novel unsupervised framework to pre-train generalist policies that capture diverse, optimal, long-horizon behaviors from unlabeled offline data such that they can be quickly adapted to any arbitrary new tasks in a zero-shot manner. Our key insight is to learn a structured representation that preserves the temporal structure of the underlying environment, and then to span this learned latent space with directional movements, which enables various zero-shot policy “prompting” schemes for downstream tasks. Through our experiments on simulated robotic locomotion and manipulation benchmarks, we show that our unsupervised policies can solve goal-conditioned and general RL tasks in a zero-shot fashion, even often outperforming prior methods designed specifically for each setting. Our code and videos are available at https://seohong.me/projects/hilp/.
@@ 3 Preliminaries and Problem Setting
Hilbert space. A Hilbert space ${\mathcal{Z}}$ is a complete vector space equipped with an inner product $\langle x,y\rangle$, the induced norm $\|x\|=\sqrt{\langle x,x\rangle}$, and the induced metric $d(x,y)=\|x-y\|$ for $x,y\in{\mathcal{Z}}$. Intuitively, a Hilbert space can roughly be thought of as a “stricter” version of a metric space, where there exists an inner product that is consistent with the metric. For example, a Euclidean space with the $\ell^{1}$- or $\ell^{\infty}$-norm is a metric space but not a Hilbert space, whereas a Euclidean space with the $\ell^{2}$-norm is a Hilbert space, as $\|x\|_{2}=\sqrt{x^{\top}x}$ for $x\in{\mathbb{R}}^{D}$. In our experiments, we will mainly employ Euclidean spaces (with the $\ell^{2}$-norm) as Hilbert spaces, but the theorems in the paper can be applied to any arbitrary real Hilbert space.
@@
Our goal is to pre-train a versatile latent-conditioned policy $\pi(a\mid s,z)$, where $z\in{\mathcal{Z}}$ denotes a latent vector (which we call a task or a skill), purely from the unlabeled offline data ${\mathcal{D}}$, without online interactions. For the evaluation of the pre-trained policy, we consider three evaluation settings. (1) Zero-shot RL [^1] We use the term “zero-shot RL” following Touati et al. (2023).: Given a reward function $r(s)$, we aim to find the best latent vector $z$ that maximizes the reward function, without additional training. (2) Offline goal-conditioned RL: Given a target goal $g\in{\mathcal{S}}$, we aim to find the best latent vector $z$ of the policy $\pi(a\mid s,z)$ that leads to the goal as quickly as possible, without additional training. The goal is specified at test time. (3) Hierarchical RL: Given a reward function $r(s)$, we train a high-level policy $\pi^{h}(z\mid s)$ that sequentially combines pre-trained skills to maximize the reward function using offline RL. In all three settings, we only allow online interaction with the environment during the final evaluation, and assume that the state space and environment dynamics remain the same at evaluation time.
@@ 4.1 Hilbert Representations
Remarks. There exist three potential limitations with Equation 6. First, the distance metric in ${\mathcal{Z}}$ is symmetric, whereas temporal distances might be asymmetric. Second, even when the environment dynamics are symmetric, there might not exist an exact isometry between the MDP and the Hilbert space (Indyk et al., 2017; Pitis et al., 2020). Third, we use a discount factor $\gamma$ in Equation 6, but temporal distances are undiscounted. In this regard, our objective might better be viewed as finding the best discounted Hilbert approximation of the MDP, rather than learning an exact Hilbert abstraction. While this might not be ideal in highly asymmetric environments, we note that we will not directly use the potentially erroneous parameterized value function $V(s,g)$ for policy learning. We will instead only take the representation $\phi$, defining a new reward function as well as a new value function to pre-train an unsupervised latent-conditioned policy, as we will describe in Section 4.2. After all, our goal is to train a representation $\phi$ that captures the long-term temporal structure of the MDP, and we empirically found that even approximate Hilbert representations lead to diverse useful behaviors that can be directly used to solve downstream tasks in our experiments (Section 6). We refer to Appendix C for further discussions.
@@ 5 Why are HILPs useful?
HILPs, in combination with structured representations $\phi(s)$, provide a number of ways to solve downstream tasks efficiently, often even in a zero-shot fashion. In this section, we describe three test-time “prompting” strategies for HILPs for zero-shot RL (Section 5.1), offline goal-conditioned RL (Section 5.2), and test-time planning (Section 5.3).
@@ 5.1 Zero-Shot RL
5.2 Offline Goal-Conditioned RL
@@ 6 Experiments
In our experiments, we empirically evaluate the performance of HILPs on various types of downstream tasks. In particular, we consider three different experimental settings: zero-shot RL, offline goal-conditioned RL, and hierarchical RL (see Section 3 for the problem statements), in which we aim to answer the following questions: (1) Can HILPs be prompted to solve goal-conditioned and reward-based tasks in a zero-shot manner? (2) Can a single general HILP framework outperform prior state-of-the-art approaches specialized in each setting? (3) Does test-time planning improve performance? In our experiments, we use $8$ random seeds unless otherwise stated, and report $95\%$ bootstrap confidence intervals in plots and standard deviations in tables. Our code is available at this repository.
@@ 6.1 Zero-Shot RL
6.2 Offline Goal-Conditioned RL
@@
Table 1: Offline goal-conditioned RL performance ($\mathbf{8}$ seeds). HILP achieves the best performance among general offline unsupervised policy learning methods, while being comparable to methods that are specifically designed to solve offline goal-conditioned RL. With our efficient test-time planning procedure based on Hilbert representations, HILPs often even outperform offline goal-conditioned RL methods.
@@
visual-kitchen-partial $\mathbf{63.2}$$\pm 3.7$ $\mathbf{63.6}$$\pm 4.2$ $43.3$$\pm 2.2$ $39.7$$\pm 3.3$ $59.9$$\pm 4.0$ $52.1$$\pm 5.3$
visual-kitchen-mixed $\mathbf{57.5}$$\pm 4.2$ $52.9$$\pm 4.7$ $50.9$$\pm 3.6$ $49.6$$\pm 5.7$ $\mathbf{55.9}$$\pm 9.7$ $54.1$$\pm 13.4$
Table 1 shows the results, which suggest HILPs can solve challenging long-horizon goal-conditioned tasks in a zero-shot manner, and significantly outperform previous general unsupervised policy learning methods (FB and three SF methods). This is likely because our method prioritizes learning long-horizon behaviors that span the state space, unlike previous successor features or forward-backward methods. Table 1 also shows that HILPs with test-time planning often even outperform GCRL-dedicated methods (e.g., GC-IQL). We note that this planning procedure can be done only with elementary algebraic operations based on pre-computed representations (Section 5.3), thanks to our structured Hilbert representations. To further study the effect of test-time planning, we compare the performances of HILPs with different numbers (00-$3$) of midpoint recursions. We report results in Figure 7, which shows that iterative refinement of prompts via our test-time planning approach improves performance consistently. We refer to Appendix B for further analyses, including an ablation study, embedding error analysis, and latent space visualization.
@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Limitations
As mentioned in Section 4.1, one limitation of our Hilbert representation objective is that a symmetric Hilbert space might not be expressive enough to capture arbitrary MDPs. Although we empirically demonstrate that HILPs exhibit strong performances in various complex, long-horizon simulated robotic environments, they might struggle in highly asymmetric or disconnected MDPs (e.g., environments in which gravity plays a significant role), where there might not exist a reasonable approximate isometry to a Hilbert space. We believe this limitation may be resolved by combining the notion of an inner product with a universal quasimetric embedding (Pitis et al., 2020; Wang et al., 2023), which we leave for future work. Another limitation is that our value-based representation learning objective (Equation 6) might be optimistically biased in stochastic or partially observable environments (Park et al., 2023). We believe combining our method with history-conditioned policies or recurrent world models may be one possible solution to deal with such MDPs. We also note that our experiments assume the state space and environment dynamics to be the same at evaluation time. We leave applying HILPs to multi-environment or transfer learning settings for future work. Finally, we use Euclidean spaces as Hilbert spaces for our experiments in this work. We believe applying HILPs to more general Hilbert spaces, such as the $L^{2}$ function space or reproducing kernel Hilbert spaces, may significantly enhance the expressivity of Hilbert representations.
@@
Figure 9: Ablation study of latent dimensions for offline goal-conditioned RL ($\mathbf{4}$ seeds). In D4RL tasks, $D=32$ generally leads to the best performance across the environments.
Figure 10: Embedding error analysis ($\mathbf{4}$ seeds). We show the relationship between performance and Hilbert embedding errors on the antmaze-large-diverse task. In general, lower embedding errors lead to better goal-reaching performance.
@@ Appendix B Additional Results
Ablation study. Figures 9 and 8 show how the dimension $D$ of the latent space ${\mathcal{Z}}={\mathbb{R}}^{D}$ affects the performances of zero-shot RL and offline goal-conditioned RL (without planning), where we use $D=50$ for zero-shot RL and $D=32$ for goal-conditioned RL in our main experiments. The results suggest that a latent dimension between $25$ and $64$ generally leads to the best performance in both cases.
Embedding error analysis. To understand the relationship between Hilbert embedding errors and goal-reaching performance, we compare the mean squared errors (MSEs) of Hilbert representations (i.e., $\mathbb{E}[(d^{*}(s,g)-\|\phi(s)-\phi(g)\|)^{2}]$) and the final performances with different embedding dimensions on the antmaze-large-diverse task. To approximate the ground-truth temporal distance $d^{*}(s,g)$ in practice, we employ a monolithic goal-conditioned value function $V(s,g)\approx-d^{*}(s,g)$ trained with a separate goal-conditioned IQL objective. We use the same IQL expectile of $0.9$ for both value functions in this analysis. Figure 10 shows the results, suggesting that low embedding errors generally lead to better goal-reaching performances, as predicted by Theorem 5.1.
@@
Full results on the ExORL benchmark. Tables 3 and 4 present the full results (unnormalized returns) on the state- and pixel-based ExORL benchmarks. Figures 12, 14, 13, 15, 17 and 16 show plots with three different aggregation criteria (per-environment, per-dataset, and per-task) on the state- and pixel-based ExORL benchmarks.
@@ Appendix C Theoretical Results
Limitations. One natural question to ask is whether it is always possible to embed any MDP into a Hilbert space up to arbitrary accuracy. Unfortunately, this is not always possible. First, temporal distances are asymmetric but the distance metric of the Hilbert space ${\mathcal{Z}}$ is symmetric. Second, even when the environment is completely symmetric, there exists a symmetric MDP that is not embeddable into a Hilbert space with an arbitrarily low approximation error (Indyk et al., 2017; Pitis et al., 2020). This is mainly because Hilbert spaces are highly structured, especially compared to metric or quasimetric spaces, which do not require a well-defined inner product. Nonetheless, the inner product structure of the Hilbert space naturally enables useful prompting strategies for zero-shot RL and goal-conditioned RL, and we empirically found that even MDPs that in principle do not have a lossless Hilbert representation can still be solved effectively via our method in our experiments. We believe finding a way to relax the Hilbert condition while having similar prompting and planning strategies is an exciting and important future research direction.
@@ Appendix D Experimental Details
We implement HILPs based on two different codebases: the official implementation of FB representations (Touati et al., 2023) for zero-shot RL experiments and that of HIQL (Park et al., 2023) for offline goal-conditioned RL and hierarchical RL experiments. Our implementations are publicly available at the following repository: https://github.com/seohongpark/HILP. We run our experiments on an internal cluster consisting of A5000 GPUs. Each run in this work takes no more than $28$ hours.
@@ D.2 Implementation Details
Hilbert representations. In Equation 6, we use the same goal relabeling strategy as Park et al. (2023) except that we do not set $g=s$, since $V(s,s)=0$ is always guaranteed in our parameterization. Namely, we sample $g$ either from a geometric distribution over the future states within the same trajectory (with probability $0.625$), or uniformly from the dataset (with probability $0.375$). We note that the values $0.625$ and $0.375$ come from the original hyperparameters used by Park et al. (2023) (which use $g=s$ with probability $0.2$, future states with probability $0.5$, and random states with probability $0.3$), where we redistribute the unnecessary probability mass of $g=s$ across the other two bins. To avoid numerical issues with gradient descent, we add a small value (${\varepsilon}=10^{-6}$) when computing $\|\phi(s)-\phi(g)\|$.
@@
Offline goal-conditioned RL (Section 6.2). We implement HILP on top of the official codebase of the work by Park et al. (2023). For HILP, we use the reward function in Equation 7 and the goal-conditioned prompting scheme introduced in Section 5.2. We use IQL (Kostrikov et al., 2022) with AWR (Peng et al., 2019) as an offline algorithm to train policies. For HILP-Plan, at each evaluation epoch, we first randomly sample $N=50000$ states $w_{1},w_{2},\dots,w_{N}$ from the dataset ${\mathcal{D}}$, and pre-compute their representations $\phi(w_{1}),\phi(w_{2}),\dots,\phi(w_{N})$. Then, at every time step, we find the $\operatorname*{arg\,min}$ of Equation 10 over the $N$ samples using the pre-computed representations. In practice, we use the average of the $50$ $\operatorname*{arg\,min}$ representations, as we found this to lead to better performance. For GC-IQL and GC-BC, we use the implementations provided by Park et al. (2023). They are implemented on the same codebase as HILP. For GC-CQL, we modify the JaxCQL repository (Geng, 2022) to make it compatible with our goal-conditioned setting. We mostly follow the hyperparameters used by Nakamoto et al. (2023). We use the same goal relabeling strategy as Park et al. (2023) for all three goal-conditioned RL methods. For FB, we use the official implementation provided by Touati et al. (2023), where we additionally implement D4RL environments. For SF methods (FDM, Lap, and Rand), we re-implement IQL versions of them on the same codebase as HILP, as we found these versions to perform better than the original implementations by Touati et al. (2023). Among FB and SF methods, we only re-implement SF methods based on IQL, as FB in its current form is not directly compatible with IQL. We report the full list of the hyperparameters used in our offline goal-conditioned RL experiments in Table 6.
@@
MLP dimensions $(512,512)$ ($\phi$), $(1024,1024,1024)$ (others)222 Following Touati et al. (2023), for policies $\pi(a\mid s,z)$ and TD3 values $Q(s,a,z)$, we process $s$ (or $(s,a)$) and $(s,z)$ separately with $(1024,512)$-sized MLPs, concatenate them together, and then pass another $(1024)$-sized MLP.
@@
Table 6: Hyperparameters for offline goal-conditioned RL.
@@
Policy MLP dimensions $(256,256)$ ({GC-BC, GC-IQL} on AntMaze-Large)333 We found that $(256,256)$-sized policy networks lead to better performance than $(512,512,512)$-sized ones for GC-BC and GC-IQL on AntMaze-Large. , $(512,512,512)$ (others)
@@
Latent dimension $32$ (HILP), $8$ (OPAL)444 We found that OPAL works better with $8$-dimensional latent spaces (as in the original work), compared to $32$-dimensional ones (as in this work), especially in AntMaze tasks.